# Transfer learning via Regularized Linear Discriminant Analysis

### v2

@@ 3.2 Pooled Sample Covariance and Individual Sample Covariance Matrix
For TLP-RDA, all discriminant directions uses the same covariance estimate $\widehat{\Sigma}_{P}$. Although $\widehat{\Sigma}_{P}$ is a better covariance estimate than all $\widehat{\Sigma}_{k}$, $\widehat{d}_{k}^{P}$ are inevitably similar and the columns space of $\widehat{d}_{k}^{P}$ would be less informational. This essentially becomes a bias-variance trade off. When $\gamma_{k}$ is small, the estimates $\widehat{d}_{k}$ are reliable already. In this case, individual covariance matrices bring more variances to this column space, therefore, increase the quality of final TL-RDA estimator. When $\gamma_{k}$ are large, direction estimate based on $\widehat{\Sigma}_{k}$ are no longer trustworthy and one should consider the more stable $\widehat{\Sigma}_{P}$. We will formalize this statement in this section also.
@@
We also numerically demonstrate this phenomenon by plugging values into the limiting expressions under a more general set up (figure 3). We again use $\rho=0.5$ while changing the other parameters, including the number of auxiliary studies $K$, the decay rate of eigenvalues of $\Sigma$ and the signal strength $\alpha$. We can see TLP-RDA is able to outperform TL-RDA in all cases as $\gamma_{K}$ grows larger, as the error rates of TLP-RDA decrease at slower rates. The transition point $\gamma^{*}$, defined as the $\gamma_{K}$ when TLP-RDA outperforms TL-RDA, differs in different scenarios. One can see it decreases when the eigenvalues of $\Sigma$ decreases slower, and interestingly, also when the number of auxiliary population decreases.
@@ 4 Heterogeneous Population Covariance Matrix
Note that this is identical to the original set up assumption 1.1 except each population has a different population covariance matrix $\Sigma_{k}$. In addition, we assume the assumption 1.2 holds for all covariance matrices $\Sigma_{k}$.
@@
We can then derive the optimal prediction and estimation weights in a manner identical to Theorems 2.3 and 2.7. Note that the Bayes optimal discriminant direction is given by $\Sigma_{K}^{-1}\delta_{K}$, and as before we aim to leverage the related observations in each source through a weighted linear combination of their discriminant directions $\widehat{d}_{k}$, for $k=1,\dots,K$.
@@
Suppose that assumptions 1.3, 1.4, 2.1, 2.2 as well as 4.1 and 4.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the error in estimating the Bayes optimal discriminator $d_{Bayes}$ is given by:
@@
Suppose that assumptions 1.3, 1.4 as well as 4.1 and 4.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the excess risk, i.e., the error in predicting the class at a random test point $x_{0}$, when compared to the Bayes optimal discriminator $d_{Bayes}$, is given by:
@@
In order to utilize the above weights in a practical scenario, we need to replace all unknown quantities by their estimates. As before, the Marcenko Pastur law is crucial in estimating the parameters related to the spectral distributions of $\Sigma_{k}$, for $k=1,\dots,K$. In addition, we take a different approach from Tripuraneni et al. (2021); Song et al. (2024) to avoid directly characterizing the joint spectral distribution of covariance matrices in different populations. Therefore, our estimations of optimal weights do not depend on unknown constants nor require additional assumptions. Most of the details are straightforward and identical in estimation procedure as the rest of the paper. It remains to provide consistent estimators for ${\cal Y}_{kk^{\prime}}$, which we describe in the following proposition.
@@
5 Real data
We utilize the CRIC dataset to evaluate the performance of the proposed transfer learning methods. The dataset comprises observations from $2,182$ subjects, with $4,830$ protein measurements collected. These subjects were approximately evenly distributed across seven university sites. We hypothesize that while the mechanisms by which proteins predict events are related across these sites, they are not identical. Consequently, we sequentially designate each site as the target population, with the remaining sites serving as auxiliary populations, aiming to enhance prediction accuracy using TL-RDA or TLP-RDA.
@@ 5 Real data
Table 1: AUC on testing data. The TL column show the testing AUC of the transfer learning RDA method with highest cross-validated AUC. The TL type columns show the type of the selected transfer learning RDA method. TL-RDA-E, TL-RDA-P, TLP-RDA-E, TLP-RDA-P are respectively TL-RDA with the optimal estimation / prediction weights, TLP-RDA with the optimal estimation / prediction weights.
@@
The bold numbers in each row of the table indicate the best-performing algorithm, and the transfer learning methods that lose to the competing methods are colored red. It is evident that the transfer learning RDA methods outperform the others in the vast majority of cases. This demonstrates that not only are TL-RDAs capable of borrowing information to improve prediction accuracy, but they also do so more efficiently than simply pooling the data. Notably, TLP-RDA with optimal prediction weights outperforms the other methods most frequently. The amount of improvement brough by transfer learning also seemingly increase as we include more covariates. Additionally, we observe that methods based on pooled sample covariance tend to perform better when the number of features $p$ is larger, consistent with Proposition 3.2.
@@ A.2 Technical Lemmas and Their Proofs
Under assumptions 1.2-1.4, consider a matrix $\mathbf{Z}\in\mathbb{R}^{p\times p}$ whose entries $Z_{ij}$ have finite $(8+\epsilon)$-th moment for some $\epsilon>0$. Suppose further that $\mathbf{Z}$ is independent of $\delta_{k}$, $k,k^{\prime}=1,\cdots,K;k\neq k^{\prime}$, we have as $p\rightarrow\infty$
@@
(3). Under the Assumptions 2.1 and 2.2, we have
@@ A.3 Proofs of Theorems
Under Assumptions 1.3 and 1.4; by Lemma 3.7 in Dobriban and Wager (2018), we know $\widehat{b}\rightarrow_{a.s.}0$. By Lemma 3.8 in the same paper, we can use the almost sure limit $(\widehat{d}(\mathbf{w}))^{\top}\mu_{-1}\rightarrow_{a.s.}(\widehat{d}(\mathbf{w}))^{\top}\delta_{K}$ and arrive at
@@ B.1 Validation of Limiting Error Formulae
In this subsection, we use synthetic data to verify the accuracy of the theoretical formula in Theorems 2.1, 2.3, 2.7. We use a Toeplitz-type covariance matrix, $n_{1},\cdots,n_{K}=150,140,130,120,110,100$ and $p=150$. We use the same signal strength for all population $\alpha^{2}=0.5$ and the same weight correlation $\rho=0.5$. The prediction error is based on $2000$ simulated testing data points. We then vary $\lambda_{k}$ from $0.3$ to $10$. For each $\lambda_{k}$, the simulation is repeated $50$ times. One can refer to figure 4 and figure 5 for the comparison between simulated testing error rate (boxplots) and theoretical error rate (red line) for TL-RDA and TL-RDA respectively. We can see that the limiting formulae are very accurate across the $\lambda_{k}$ considered.
@@ Appendix C Limiting Error and Optimal Weight under Unequal Sampling
Recall the error rate under Assumption 1.1 is

### v3

@@ 5.2 Pooled Sample Covariance and Individual Sample Covariance Matrix
For TLP-RDA, all discriminant directions uses the same covariance estimate $\widehat{\Sigma}_{P}$. Although $\widehat{\Sigma}_{P}$ is a better covariance estimate than all $\widehat{\Sigma}_{k}$, $\widehat{d}_{k}^{P}$ are inevitably similar and the columns space of $\widehat{d}_{k}^{P}$ would be less informational. This essentially becomes a bias-variance trade off. When $\gamma_{k}$ is small, the estimates $\widehat{d}_{k}$ are reliable already. In this case, individual covariance matrices bring more variances to this column space, therefore, increase the quality of final TL-RDA estimator. When $\gamma_{k}$ are large, direction estimate based on $\widehat{\Sigma}_{k}$ are no longer reliable and one should consider the more stable $\widehat{\Sigma}_{P}$. We will formalize this statement in this section also.
@@
We also numerically demonstrate this phenomenon by plugging values into the limiting expressions under a more general set up (Figure 3). We again use $\rho=0.5$ while changing the other parameters, including the number of auxiliary studies $K$, the decay rate of the eigenvalues of $\Sigma$ and the signal strength $\alpha$. We can see that TLP-RDA can outperform TL-RDA in all cases as $\gamma_{K}$ grows, as the error rates of TLP-RDA decrease at slower rates. The transition point $\gamma^{*}$, defined as the $\gamma_{K}$ when TLP-RDA outperforms TL-RDA, differs in different scenarios. One can see that it decreases when the eigenvalues of $\Sigma$ decrease slower, and interestingly, also when the number of auxiliary population decreases.
@@ 6 Heterogeneous Population Covariance Matrix
Note that this is identical to the original set up except each population has a different population covariance matrix $\Sigma_{k}$. In addition, we assume the assumption 3.3 holds for all covariance matrices $\Sigma_{k}$.
@@
We can then derive the optimal prediction and estimation weights in a manner identical to Theorems 4.3 and 4.7. Since the Bayes optimal discriminant direction is given by $\Sigma_{K}^{-1}\delta_{K}$, and as before we aim to leverage the related observations in each source through a weighted linear combination of their discriminant directions $\widehat{d}_{k}$, for $k=1,\dots,K$.
@@
Suppose that assumptions 3.1, 3.2, 4.1, 4.2 as well as 6.1 and 6.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the error in estimating the Bayes optimal discriminator $d_{Bayes}$ is given by:
@@
Suppose that assumptions 3.1, 3.2 as well as 6.1 and 6.2 hold. Then for a fixed $K\geq 2$, as $n_{k},p\rightarrow\infty,p/n_{k}\rightarrow\gamma_{k}\in(0,\infty]$ for $1\leq k\leq K$, the weight for minimizing the excess risk, i.e., the error in predicting the class at a random test point $x_{0}$, when compared to the Bayes optimal discriminator $d_{Bayes}$, is given by:
@@
In order to utilize the above weights in a practical scenario, we need to replace all unknown quantities by their estimates. As before, the Marcenko Pastur law is crucial in estimating the parameters related to the spectral distributions of $\Sigma_{k}$, for $k=1,\dots,K$. Most of the details are straightforward and identical in estimation procedure as the rest of the paper. It remains to provide consistent estimators for ${\cal Y}_{kk^{\prime}}$, which we describe in the following proposition.
@@
7 Proteomics-based Prediction of 10-year Cardiovascular Disease Risk
We utilize the data set from the Chronic Renal Insufficiency Cohort (CRIC) study to evaluate the performance of the proposed transfer learning methods. The dataset comprises observations from $2,182$ subjects who had not had any cardiovascular disease at the baseline, with $4,830$ protein measurements collected. These subjects were approximately evenly distributed across seven university sites. Our goal is to build a classification/prediction model for 10-year cardiovascular disease risk based on the baseline plasma proteomic data. We hypothesize that while the mechanisms by which proteins predict events are related across these sites, they are not identical. Consequently, we sequentially designate each site as the target population, with the remaining sites serving as auxiliary populations, aiming to enhance prediction accuracy using TL-RDA or TLP-RDA.
@@ 7 Proteomics-based Prediction of 10-year Cardiovascular Disease Risk
Table 1: AUC on testing data for 10-year cardiovascular disease risk in CRIC cohort. The TL column show the testing AUC of the transfer learning RDA method with highest cross-validated AUC. The TL type columns show the type of the selected transfer learning RDA method. TL-RDA-E, TL-RDA-P, TLP-RDA-E and TLP-RDA-P are respectively TL-RDA with the optimal estimation / prediction weights, TLP-RDA with the optimal estimation / prediction weights. The bold numbers in each row of the table highlight the best-performing algorithm. Transfer learning methods that perform worse than the competing methods are highlighted in red.
@@
The bold numbers in each row of the table highlight the best-performing algorithm. Transfer learning methods that perform worse than the competing methods are highlighted in red. It is evident that the transfer learning RDA methods outperform the others in the vast majority of cases. This demonstrates that not only are TL-RDAs capable of borrowing information to improve prediction accuracy, but they also do so more efficiently than simply pooling the data. Notably, TLP-RDA with optimal prediction weights outperforms the other methods most frequently. The amount of improvement brought by transfer learning also seemingly increase as we include more proteins in the model. Additionally, we observe that methods based on pooled sample covariance tend to perform better when the number of features $p$ is larger, consistent with Proposition 5.2.
8 Discussion
We have developed methods of transfer learning for classification using regularized random effects linear discriminant analysis. In this approach, the discriminant direction is estimated through a weighted combination of the regularized estimates of discriminant directions derived from both the target and source models. By leveraging results from random matrix theory, we have demonstrated that the weights, as well as the classification error rate, can be accurately estimated using the data. Our findings highlight that the optimal choice of weights depends critically on the underlying true models and the presence or absence of distributional shifts between the testing and training data in the target model.
Through comprehensive empirical evaluations, we have illustrated the practical utility of the proposed methods. Specifically, we applied the transfer learning approach to predict 10-year cardiovascular disease risk using high-dimensional protein expression data. The results demonstrate that incorporating information from related source datasets substantially improved classification performance compared to target-only models or models based on pooled data. This improvement underscores the importance of accounting for shared information across datasets while accommodating potential distributional differences.
This paper focuses on two-class LDA models. A promising direction for future research is extending these methods to multi-class classification within the framework of transfer learning. Another interesting avenue is exploring this LDA problem under privacy or communication constraints, as has been recently studied for other classification methods (see, e.g., Auddy et al., 2024).
Acknowledgments
This research was supported by NIH grants R01GM129781 and U01DK108809.
@@ A.2 Technical Lemmas and Their Proofs
Under assumptions 3.3-3.2, consider a matrix $\mathbf{Z}\in\mathbb{R}^{p\times p}$ whose entries $Z_{ij}$ have finite $(8+\epsilon)$-th moment for some $\epsilon>0$. Suppose further that $\mathbf{Z}$ is independent of $\delta_{k}$, $k,k^{\prime}=1,\cdots,K;k\neq k^{\prime}$, we have as $p\rightarrow\infty$
@@
(3). Under the Assumptions 4.1 and 4.2, we have
@@ A.3 Proofs of Theorems
Under Assumptions 3.1 and 3.2; by Lemma 3.7 in Dobriban and Wager (2018), we know $\widehat{b}\rightarrow_{a.s.}0$. By Lemma 3.8 in the same paper, we can use the almost sure limit $(\widehat{d}(\mathbf{w}))^{\top}\mu_{-1}\rightarrow_{a.s.}(\widehat{d}(\mathbf{w}))^{\top}\delta_{K}$ and arrive at
@@ B.1 Validation of Limiting Error Formulae
In this subsection, we use synthetic data to verify the accuracy of the theoretical formula in Theorems 4.1, 4.3, 4.7. We use a Toeplitz-type covariance matrix, $n_{1},\cdots,n_{K}=150,140,130,120,110,100$ and $p=150$. We use the same signal strength for all population $\alpha^{2}=0.5$ and the same weight correlation $\rho=0.5$. The prediction error is based on $2000$ simulated testing data points. We then vary $\lambda_{k}$ from $0.3$ to $10$. For each $\lambda_{k}$, the simulation is repeated $50$ times. One can refer to figure 4 and figure 5 for the comparison between simulated testing error rate (boxplots) and theoretical error rate (red line) for TL-RDA and TL-RDA respectively. We can see that the limiting formulae are very accurate across the $\lambda_{k}$ considered.
@@ Appendix C Limiting Error and Optimal Weight under Unequal Sampling
Recall the error rate under Assumption 1 is